Use Slater's rules to estimate Zeff for the nickel, copper, zinc and gallium atom
Answer #1

US ① Mi => 152 25² 2pt 35² 3pc 348 45² (1X0-35 ) + ( 16x0.85) + (10x1) (n-2 6- = 0.35+ 13.6+ 10 = 23.95 Zell 7-6, 28- 23.95 =thank you vmuch....